### v1

# Surprisal from Larger Transformer-based Language Models Predicts fMRI Data More Poorly

Abstract
As Transformers become more widely incorporated into natural language processing tasks, there has been considerable interest in using surprisal from these models as predictors of human sentence processing difficulty. Recent work has observed a positive relationship between Transformer-based models’ perplexity and the predictive power of their surprisal estimates on reading times, showing that language models with more parameters and trained on more data are less predictive of human reading times. However, these studies focus on predicting latency-based measures (i.e., self-paced reading times and eye-gaze durations) with surprisal estimates from Transformer-based language models. This trend has not been tested on brain imaging data. This study therefore evaluates the predictive power of surprisal estimates from 17 pre-trained Transformer-based models across three different language families on two functional magnetic resonance imaging datasets. Results show that the positive relationship between model perplexity and model fit still obtains, suggesting that this trend is not specific to latency-based measures and can be generalized to neural measures.
Surprisal from Larger Transformer-based Language Models
Predicts fMRI Data More Poorly
Yi-Chien Lin The Ohio State University Department of Linguistics lin.4434@osu.edu                        William Schuler The Ohio State University Department of Linguistics schuler.77@osu.edu
1 Introduction
Expectation-based theories of syntactic comprehension (Hale, 2001; Levy, 2008) posit that comprehenders evaluate multiple hypotheses of a structure in parallel. Within this line of research, the processing difficulty of a word is assumed to be inversely proportional to the predictability of that word given the context which has been exposed to the comprehender. Processing difficulty can be quantified using information-theoretic surprisal (Shannon, 1948) and the processing difficulty of a word is proportional to its surprisal (negative log probability). As Transformer-based models (Vaswani et al., 2017) become more widely incorporated into natural language processing (NLP) tasks, there has been considerable interest in using surprisal from these models as predictors of processing difficulty (Merkx and Frank, 2021; Wilcox et al., 2020).
Recent work has observed that surprisal from larger Transformer-based language models (LMs), which have more parameters and lower model perplexity, is less predictive of reading times (Oh et al., 2022; Oh and Schuler, 2023; de Varda and Marelli, 2023; Shain et al., 2024; Oh et al., 2024). Larger LMs trained on larger amounts of data become more predictive of rare tokens throughout the course of training, resulting in lower surprisal estimates that diverge from human reading times (Oh et al., 2024). In addition to reading times, Oh et al. (2022) also use surprisal estimates from Transformer-based models to predict brain imaging data, but observe no inverse scaling using a relatively small set of LMs. However, moving away from surprisal, recent studies observed a positive scaling on brain imaging data when using vectors directly from large language models as predictors (Schrimpf et al., 2021; Hosseini et al., 2024) so it may be tempting to think brain imaging data behaves differently from latency-based data.
This study therefore bridges the gap by evaluating the inverse scaling trend of the predictive power of surprisal estimates from different Transformer-based model variants on brain imaging data. The goal is to examine whether the positive relationship between model perplexity and the predictive power of surprisal estimates still holds for brain imaging data. Results show that the positive relationship still obtains when evaluating the predictive power of surprisal estimates on brain imaging data, suggesting that this trend is not specific to latency-based measures.
2 Related Work
Previous studies have examined the predictive power of surprisal estimates from various types of LMs such as n-gram and recurrent neural network models on psychometric data (e.g., Goodkind and Bicknell, 2018; Aurnhammer and Frank, 2019; Wilcox et al., 2020). These studies observed a consistent trend in the positive relationship between the a model’s quality and its predictive power on psychometric data such as reading times: the better an LM’s quality (or the lower an LM’s perplexity), the better its fit to the psychometric data. Goodkind and Bicknell (2018) compared the predictive power of surprisal estimates from n-gram and Long Short-Term Memory network (LSTM; Hochreiter and Schmidhuber, 1997) models and found that a model’s perplexity and its ability to predict reading times are linearly and negatively correlated. Similarly, Wilcox et al. (2020) observed a negative relationship between model perplexity and model fit when evaluating surprisal estimates from LSTMs, recurrent neural network grammars (Dyer et al., 2016), Transformers, and n-gram models.
Recently, due to the dominance of Transformer-based models in the field of NLP, there has been increasing interest in comparing surprisal estimates from these models against behavioral and neural measures of processing difficulty. Differing from previous findings, Oh et al. (2022) observed a completely opposite positive relationship between LM perplexity and models’ fit to reading times when using surprisal estimates from a set of pre-trained GPT-2 models (Radford et al., 2019). In other words, larger LMs (with lower model perplexity) are less predictive of reading times. This observation was further replicated using surprisal estimates from a larger set of Transformer-based LMs (Oh and Schuler, 2023; Oh et al., 2024).
Moreover, recent work in this line of study found that surprisal estimates from neural LMs have a tendency to underpredict reading times. Van Schijndel and Linzen (2021) and Arehalli et al. (2022) showed that neural LM surprisal successfully predicts human precessing difficulty in garden-path constructions but consistently underestimate its magnitude. Kuribayashi et al. (2022) also observed that surprisal estimates from neural LMs without context limitations underpredict reading times of English and Japanese naturalistic text. Furthermore, Oh and Schuler (2023) conducted an analysis of the positive relationship between Transformer-based LM perplexity and the predictive power of surprisal estimates, observing that the poorer fit to reading times of larger LMs is mainly driven by the lower surprisal values those LMs assign to open-class words such as nouns and adjectives. Oh et al. (2024) followed up on this and showed that word frequency explains the inverse correlation between the size of Transformer-based LMs, the amount of training data, and the predictive power of surprisal estimates. They found that larger LMs get better at predicting rare words, resulting in lower surprisal.
3 Experimental Setup
This experiment collects surprisal estimates from model variants of three autoregressive Transformer-based LM families, GPT-2 (Radford et al., 2019), GPT-Neo (Black et al., 2021, 2022; Wang and Komatsuzaki, 2021), and OPT (Zhang et al., 2022). Surprisal estimates are then used to predict blood oxygenation level-dependent (BOLD) signals from two functional magnetic resonance imaging (fMRI) datasets, Shain et al. (2020) (hereafter ‘Natural Stories fMRI’) and Pereira et al. (2018) (hereafter ‘Pereira fMRI’). The experiment setup followed that of Shain et al. (2020).
3.1 Response Data
Natural Stories fMRI (Shain et al., 2020) was collected at a two-second fixed time interval from 78 subjects while they listen to a recording of the Natural Stories Corpus (Futrell et al., 2021) which consists of naturalistic English stories. The fMRI data includes the time series of BOLD signals of several functional regions of interest collected with fMRI.
Pereira fMRI contains data from Experiments 2 and 3 in Pereira et al. (2018). Eight subjects participated in Experiment 2 and read 96 English passages, which includes 384 sentences in total; six subjects participated in Experiment 3 and read 72 English passages, which consists of 243 sentences in total. Five subjects participated in both experiments. Each passage consists of three to four sentences and sentences were presented to subjects one at a time for four seconds, followed by a four-second interval. Each passage was presented three times to the subject. The responses of this dataset are per-sentence BOLD signals, which were obtained by aggregating activation across regions.
Following Shain et al. (2020), both datasets were partitioned into fit, exploratory, and held-set sets which contain around 50%, 25%, and 25% of the data points. For Natural Stories fMRI, this resulted in 100,084, 51,393, and 50,818 data points in the fit, exploratory, and held-out sets. For Pereira fMRI, this resulted in the sizes of 2,268, 1,132, and 1,130 data points for fit, exploratory, and held-out sets. The regression models were fit using the fit partition and the results reported in this study are on the held-out partition.
3.2 Predictors
Surprisal estimates are defined as the negative log probabilities of words given their preceding context. We collected surprisal estimates from model variants of three autoregressive Transformer-based LM families which were used in Oh et al. (2022) and Oh and Schuler (2023): GPT-2 (Radford et al., 2019), GPT-Neo (Black et al., 2021, 2022; Wang and Komatsuzaki, 2021), and OPT (Zhang et al., 2022) [^1] Due to the constraints in computational resources available to us, we did not include the largest OPT model variant, which has around 175 billion parameters. families. The hyperparameters of each model evaluated in this study are detailed in Table 1.
Each story of the Natural Stories Corpus and passage of Pereira fMRI was first tokenized using the byte-pair encoding (Sennrich et al., 2016) tokenizer corresponding to each model variant. The tokenized texts were then input to each LM to calculate token-wise surprisal estimates. In cases where a tokenized story or passage exceeds a single context window of the LMs, to calculate the surprisal estimates for the remaining tokens, the latter half of the previous context window was used as the first half of the context window.
Recent studies (Oh and Schuler, 2024; Pimentel and Meister, 2024) have shown that, for languages which use whitespace characters in their orthography, the leading whitespaces (i.e., the whitespace characters which immediately precede the tokens) added by LMs’ tokenizers can result in a sum over all word probabilities that is greater than one. This study therefore incorporated whitespace-trailing decoding into surprisal calculation. For each word, the probability of the leading whitespace was re-allocated to its preceding word when calculating word-level probabilities.
Since the response data of Natural Stories fMRI is the time series of BOLD signals, we first collected the surprisal estimates of the words and then applied a hemodynamic response function to convolve those surprisal estimates. Since the response data of Pereira fMRI is per-sentence BOLD signals, we used the surprisal estimate of the sentence-final word as the predictor for each response data point. The motivation of using the sentence-final word surprisal is that its preceding context has already been taken into consideration when calculating the surprisal estimates of the last words of the sentences.
Model Variant #L #H $d_{model}$ Parameters
GPT-2 Small 12 12 768 $\sim$124M
GPT-2 Medium 24 16 1024 $\sim$355M
GPT-2 Large 36 20 1280 $\sim$774M
GPT2- XL 48 25 1600 $\sim$1.6B
GPT-Neo 125M 12 12 768 $\sim$125M
GPT-Neo 1.3B 24 16 2048 $\sim$1.3B
GPT-Neo 2.7B 32 20 2560 $\sim$2.7B
GPT-J 6B 28 16 4096 $\sim$6B
GPT-NeoX 20B 44 64 6144 $\sim$20B
OPT 125M 12 12 768 $\sim$125M
OPT 350M 24 16 1024 $\sim$350M
OPT 1.3B 24 32 2048 $\sim$1.3B
OPT 2.7B 32 32 2560 $\sim$2.7B
OPT 6.7B 32 32 4096 $\sim$6.7B
OPT 13B 40 40 5120 $\sim$13B
OPT 30B 48 56 7168 $\sim$30B
OPT 66B 64 72 9216 $\sim$66B
Table 1: Hyperparameters of the models examined in this study. #L refers to the number of layers of that model; #H refers to the number of attention heads each layer has; $d_{model}$ refers to the embedding size of the model.
3.3 Regression Modeling
Once the surprisal estimates were collected, we fit a set of linear mixed-effect (LME) models to BOLD signals with lme4 (Bates et al., 2015) using the fit partition. During regression modeling, in addition to surprisal estimates, we also included a set of baseline predictors for both corpora. For Natural Stories fMRI, the baseline predictors were the index of the word position in the sentence and the word length in characters. For Pereira fMRI, the baseline predictors were sentence length in words, index of sentence position within each passage, and index of the passage containing the sentence.
All models were fit using raw BOLD values along with the above baseline predictors. The predictors were centered and scaled before being used to fit the LME models. The following formula used for fitting the LME models for Natural Stories fMRI contains by-subject random slopes for word position (sentpos), word length (wlen), surprisal estimate (surp) of the data point, and a by-subject intercept. An additional random intercept was included for each subject.
BOLD ~ z.(sentpos) + z.(wlen) + z.(surp) +
       (1 + z.(sentpos) + z.(wlen) +
       z.(surp) | subject) + (1 | subject)
For Pereira, we fit the LME models with the following formula:
BOLD ~ z.(sentlen) + z.(passpos) + z.(lastwordsurp) +
       (1 + z.(sentlen) + z.(passpos) +
       z.(lastwordsurp) | subject) + (1 | sentid)
The LME models included by-subject random slopes for sentence length (sentlen), sentence position within passage (passpos), last-word surprisal (lastwordsurp), and a by-subject intercept. These models also included an additional random intercept for each sentence (sentid). We show the predictive power with the model log likelihood calculated on the held-out partition for both datasets.
3.4 Results
Figure 1 presents the relationship between the regression model log likelihood and log-transformed parameter count. The results of Natural Stories fMRI do not show a significant trend in either direction, replicating Oh et al. (2022), but the results on Pereira fMRI show that as parameter count (model size) increases, the regression model log likelihood decreases significantly ($p$ < 0.05 by a permutation test). [^2] Although the results of Natural Stories fMRI on the held-out partition do not show a clear relationship between predictive power and parameter count, the results of it on the exploratory partition do show a significant inverse scaling ($p$ 0.001). This suggests that for brain imaging data, the predictive power of surprisal estimates decreases as the model size increases, which is in line with previous finding on reading times (Oh and Schuler, 2023; Oh et al., 2024).
Figure 1: Predictive power of surprisal estimates from pretrained LMs on BOLD signals of Natural Stories fMRI and Pereira fMRI.
4 Discussion and Conclusion
Recent studies (Oh et al., 2022; Oh and Schuler, 2023; Oh et al., 2024) have observed a robust negative relationship between Transformer-based LMs’ size and the predictive power of surprisal estimates from those LMs on latency-based measures (i.e., self-paced reading times and eye-gaze durations). This work generalizes this result by examining the predictive power of surprisal estimates from Transformer-based models on brain imaging data. The predictive power of surprisal estimates from 17 Transformer-based LMs across three LM families was evaluated on two fMRI datasets. Results show that the inverse scaling between model size and model fit still obtains on brain imaging data, suggesting that this trend is not specific to latency-based measures and can be generalized to neural measures.
Limitations
This study attempts to examine the predictive power of surprisal estimates from Trasnformer-based models on fMRI data. LLMs evaluated in this work are trained on English text and the datasets evaluated are collected from English speakers. Therefore, these findings may or may not be replicated in other languages.
Ethics Statement
The fMRI datasets used in this study are from published work (Shain et al., 2020; Pereira et al., 2018). Details regarding the data collection and relevant procedures are described in respective publications. Since this work focuses on evaluating the predictive power of LLMs on fMRI data, the potential risks and harmful impacts posed by this study on the society appear to be minimal.