### v2

# Surprisal from Larger Transformer-based Language Models Predicts fMRI Data More Poorly

Abstract
There has been considerable interest in using surprisal from Transformer-based language models (LMs) as predictors of human sentence processing difficulty. Recent work has observed an inverse scaling relationship between Transformers’ per-word estimated probability and the predictive power of their surprisal estimates on reading times, showing that LMs with more parameters and trained on more data are less predictive of human reading times. However, these studies focused on predicting latency-based measures. Tests on brain imaging data have not shown a trend in any direction when using a relatively small set of LMs, leaving open the possibility that the inverse scaling phenomenon is constrained to latency data. This study therefore conducted a more comprehensive evaluation using surprisal estimates from 17 pre-trained LMs across three different LM families on two functional magnetic resonance imaging (fMRI) datasets. Results show that the inverse scaling relationship between models’ per-word estimated probability and model fit on both datasets still obtains, resolving the inconclusive results of previous work and indicating that this trend is not specific to latency-based measures.
Surprisal from Larger Transformer-based Language Models
Predicts fMRI Data More Poorly
Yi-Chien Lin The Ohio State University Department of Linguistics lin.4434@osu.edu          William Schuler The Ohio State University Department of Linguistics schuler.77@osu.edu
1 Introduction
Expectation-based theories of syntactic comprehension (Hale, 2001; Levy, 2008) posit that comprehenders evaluate multiple hypotheses of a structure in parallel. Within this line of research, processing difficulty can be quantified using information-theoretic surprisal (Shannon, 1948) and the processing difficulty of a word is proportional to its surprisal estimate (i.e., negative log probability given the preceding context).
As Transformer-based models (Vaswani et al., 2017) become more widely incorporated into natural language processing tasks, there has been considerable interest in using surprisal from these models as predictors of processing difficulty (Merkx and Frank, 2021; Wilcox et al., 2020). Recent work has observed that surprisal from larger Transformer-based LMs, which have more parameters and lower model perplexity, is less predictive of reading times (Oh et al., 2022; Oh and Schuler, 2023; de Varda and Marelli, 2023; Shain et al., 2024; Oh et al., 2024). Larger LMs trained on larger amounts of data become more predictive of rare tokens throughout the course of training, resulting in lower surprisal estimates that diverge from human reading times (Oh et al., 2024). In addition to reading times, Oh et al. (2022) also use surprisal estimates from Transformers to predict brain imaging data, but observe no inverse or positive scaling when using a relatively small set of LMs. However, moving away from surprisal, recent studies observed a positive scaling on brain imaging data when using vectors directly from large language models as predictors (Schrimpf et al., 2021; Hosseini et al., 2024) so it may be tempting to think brain imaging data behaves differently from latency-based data (i.e., reading times).
This study presents evidence against this possibility by evaluating the scaling trend of the predictive power of surprisal estimates from a comprehensive set of 17 Transformer-based LMs across three LM families on two different brain imaging datasets: the original dataset (Shain et al., 2020), which was also evaluated in Oh et al. (2022), and a separate dataset (Pereira et al., 2018). Results show statistically significant inverse scalings on both datasets which resolve the inconclusive results from Oh et al. (2022) on Shain et al. (2020) when using surprisal estimates from a larger set of LMs and replicate inverse scaling on Pereira et al. (2018).
This work provides an evidence showing that the inverse scaling of surprisal estimates is not specific to latency-based data. This piece of evidence provides a more complete picture of the relationship between the model sizes and predictive power of surprisal estimates on psychometric data, which could potentially be helpful in providing researchers with insights into the appropriate use of LMs in understanding human sentence processing.
2 Related Work
Previous studies have examined the predictive power of surprisal from various types of LMs such as n-gram, Long Short-Term Memory network (Hochreiter and Schmidhuber, 1997), and recurrent neural network models on psychometric data (e.g., Goodkind and Bicknell, 2018; Wilcox et al., 2020). These studies observed a consistent trend in the positive relationship between a model’s quality and its predictive power on psychometric data such as reading times: the better an LM’s quality (or the higher an LM’s per-word estimated probability), the better its fit to the psychometric data.
Recently, there has been increasing interest in comparing surprisal estimates from Transformer-based models against behavioral and neural measures of processing difficulty. Differing from previous findings, Oh et al. (2022) observed a completely opposite inverse scaling relationship between LMs’ per-word estimated probability and models’ fit to reading times when using surprisal estimates from a set of pre-trained GPT-2 models (Radford et al., 2019). In other words, larger LMs (with higher per-word estimated probability) are less predictive of reading times. This observation was further replicated using surprisal estimates from a larger set of Transformer-based LMs (Oh and Schuler, 2023; Oh et al., 2024).
Moreover, recent work in this line of study found that surprisal estimates from neural LMs have a tendency to underpredict reading times. Van Schijndel and Linzen (2021) and Arehalli et al. (2022) showed that neural LM surprisal successfully predicts human precessing difficulty in garden-path constructions but consistently underestimates its magnitude. Kuribayashi et al. (2022) also observed that surprisal estimates from neural LMs without context limitations underpredict reading times of English and Japanese naturalistic text. Furthermore, Oh and Schuler (2023) conducted an analysis of the inverse scaling relationship between Transformer-based LMs’ per-word estimated probability and the predictive power of surprisal estimates, observing that the poorer fit to reading times of larger LMs is mainly driven by the lower surprisal values those LMs assign to open-class words such as nouns and adjectives. Oh et al. (2024) followed up on this observation and showed that word frequency explains the inverse correlation between the size of Transformer-based LMs, the amount of training data, and the predictive power of surprisal estimates. They found that larger LMs get better at predicting rare words, resulting in lower surprisal which diverges from human reading times.
3 Experiment 1: Predictive Power of Surprisal on Shain et al. (2020)
Experiment 1 revisited Oh et al. (2022) and evaluated the predictive power of surprisal estimates from a larger set of Transformer-based LMs on Shain et al. (2020) (hereafter ‘Natural Stories fMRI’). We collected surprisal estimates from 17 pre-trained autoregressive Transformer-based LMs and used them to predict the blood oxygenation level-dependent (BOLD) signals from Natural Stories fMRI. The experiment setup followed that of Shain et al. (2020). [^1] Code for replicating the experiments in this study is available at https://github.com/modelblocks/modelblocks-release.
3.1 Response Data
Natural Stories fMRI (Shain et al., 2020) was collected at a two-second fixed time interval from 78 subjects while they listen to a recording of the Natural Stories Corpus (Futrell et al., 2021) which consists of naturalistic English stories. This dataset includes the time series of BOLD signals of several functional regions of interest (fROIs) in the language network (Fedorenko et al., 2011), which were identified with a localizer task. For each fROI, the BOLD signals were averaged across all voxels within that fROI.
Following Shain et al. (2020), Natural Stories fMRI was partitioned into fit, exploratory, and held-out sets which contain around 50%, 25%, and 25% of the data points, respectively. Specifically, we grouped the BOLD signals into 30-second chunks and distributed those chunks into the sets to ensure the time series data in the sets are sequential and to avoid having data which is temporally too close to each other in each set. This resulted in 100,084, 50,818, and 51,393 data points in the fit, exploratory, and held-out sets. The regression models were fit using the fit set and the results reported in this study are on the held-out set.
3.2 Predictors
We collected surprisal estimates from model variants of three autoregressive Transformer-based LM families which were used in Oh et al. (2022) and Oh and Schuler (2023): GPT-2 (Radford et al., 2019), GPT-Neo (Black et al., 2021, 2022; Wang and Komatsuzaki, 2021), and OPT (Zhang et al., 2022) [^2] Due to the constraints in computational resources available to us, we did not include the largest OPT model variant, which has around 175 billion parameters. families. The hyperparameters of each LM are detailed in Table 1, Appendix A.
Each story of the Natural Stories Corpus was first tokenized using the byte-pair encoding (Sennrich et al., 2016) tokenizer corresponding to each model variant. The tokenized texts were then input to each LM to calculate token-wise surprisal estimates. In cases where a tokenized story exceeds a single context window of the LMs, to calculate the surprisal estimates for the remaining tokens, the latter half of the previous context window was used as the first half of the context window. If a word consists of multiple subword tokens, we aggregate the token-level surprisal estimates to get the word-level surprisal estimate. [^3] Although we tokenized the text at token-level instead of word-level, the relationship between the predictive power of the surprisal and the LM quality is unlikely to make much difference, as Oh and Schuler (2025) found that there is a relatively small effect of token granularity on the relationship.
Recent studies (Oh and Schuler, 2024; Pimentel and Meister, 2024) have shown that, for languages which use whitespace characters in their orthography, the leading whitespaces (i.e., the whitespace characters which immediately precede the tokens) added by LMs’ tokenizers can result in a sum over all word probabilities that is greater than one. This study therefore incorporated whitespace-trailing decoding into surprisal calculation. For each word, the probability of the leading whitespace was re-allocated to its preceding word when calculating word-level probabilities.
Since the response data of Natural Stories fMRI is the time series of BOLD signals, we first collected the surprisal estimates of the words and then applied a hemodynamic response function (HRF; Boynton et al. 1996) to convolve those surprisal estimates so that the convolved surprisal estimates align with the response data. A canonical HRF convolves predictors into the shape of the BOLD signals when a subject is presented with a stimulus. The HRF we used for this experiment is shown in the following equation, giving magnitude $f(x)$ for offset time $x$ in seconds:
$f(x)=\frac{5.2^{5.4}x^{5.4-1}e^{-5.2x}}{\Gamma(5.4)}-0.35\,\frac{7.35^{10.8}x^{10.8-1}e^{-7.35x}}{\Gamma(10.8)}$
3.3 Regression Modeling
Once the surprisal estimates were collected, we fit a set of linear mixed-effect (LME) models to BOLD signals with lme4 (Bates et al., 2015) using the fit set. During regression modeling, in addition to surprisal estimates, we also included a set of baseline predictors, including the index of the word position in the sentence and the word length in characters. All models were fit using raw BOLD values along with the above baseline predictors. The predictors were centered, scaled, and convolved with the HRF before being used to fit the LME models. Following Barr et al. (2013), during model fitting, we started with the maximal LME models and eliminated one random effect at a time until the LME models converge. The maximal converging formula for Natural Stories fMRI was:
BOLD ~ z.(sentpos) + z.(wlen) + z.(surp) +
       (1 + z.(wlen) + z.(surp) | subject) +
       (1 + z.(surp) | storyid)
where sentpos is the word position in the sentence, wlen is word length, surp is surprisal estimate, and storyid is the index of a story. [^4] The original design of the LME model for Natural Stories fMRI is shown in Appendix B. We show the predictive power with the model log likelihood calculated on the held-out set.
Figure 1: Predictive power of surprisal estimates from pretrained LMs on BOLD signals of Natural Stories fMRI (left) and Pereira fMRI (right).
3.4 Results
Figure 1 (left) presents the relationship between the regression model log likelihood and log-transformed parameter count for Natural Stories fMRI. The results show a significant decrease in predictive power when parameter count (model size) increases ($p$ < 0.01 by a permutation test with 1,000 permutations), resolving the inconclusive results in Oh et al. (2022), which only evaluate the predictive power of surprisal estimates from a relatively small set of LMs on this same dataset. [^5] The choice of model families was based on previous work which observed inverse scaling on reading times (Oh et al., 2022; Oh and Schuler, 2023). Based on the inverse scaling trend observed for Pythia and the similarity to inverse scaling results for larger off-the-shelf LLMs found in Oh and Schuler (2023); Oh et al. (2024), we expect more recent, larger models to yield substantially worse fits.
4 Experiment 2: Predictive Power of Surprisal on Pereira et al. (2018)
Experiment 1 resolves the inconclusive results on Natural Stories fMRI from Oh et al. (2022) by showing an inverse scaling phenomenon with a larger set of LMs. To examine whether this phenomenon is generalizable across other brain imaging datasets, we additionally evaluated the predictive power of the same set of LMs on a different fMRI dataset, Pereira et al. (2018) (hereafter ‘Pereira fMRI’).
4.1 Response Data
Pereira fMRI contains data from Experiments II and III in Pereira et al. (2018). Eight subjects participated in Experiment II and read 96 English passages, which includes 384 sentences in total; six subjects participated in Experiment III and read 72 English passages, which consists of 243 sentences in total. Each passage consists of three to four sentences and sentences were presented to subjects one at a time. Like Natural Stories fMRI, Pereira fMRI also includes BOLD signals from the voxels in the language network. The responses of this dataset are per-sentence BOLD signals, which were obtained by aggregating activation across regions. Following Oh et al. (2022), Pereira fMRI was partitioned into fit, exploratory, and held-out sets, resulting in the sizes of 2,268, 1,132, and 1,130 data points respectively in each set.
4.2 Procedure
In general, the procedure of Experiment 2 followed that of Experiment 1 except that since the response data for Pereira fMRI is per-sentence BOLD signals, we followed recent work on the same dataset (Hosseini et al., 2024) and used the surprisal estimate of the sentence-final word as the predictor for each response data point. [^6] Hosseini et al. (2024) use the model representations of the sentence-final words and not the surprisal estimates as the representation of the entire sentence. The motivation of using the sentence-final word surprisal is that its preceding context has already been taken into consideration when calculating the surprisal estimates of the last words of the sentences.
When fitting the LME models, we used sentence length in words and index of the passage containing the sentence as baseline predictors. Similar to the procedure in Experiment 1, following Barr et al. (2013), we started with the maximal LME models and eliminated the random effects one by one until the models no longer have serious convergence issues. The maximal converging formula was:
BOLD ~ z.(sentlen) + z.(passpos)+ z.(lastwordsurp) +
       (1 + z.(sentlen) + z.(passpos) +
        z.(lastwordsurp) | subject) + (1 + z.(sentlen) +
        z.(lastwordsurp) | sentid)
where sentlen is sentence length in words, passpos is the sentence position within passage, lastwordsurp is last-word surprisal estimate, and sentid is the index of each sentence. [^7] The original design of the LME model for Pereira fMRI is shown in Appendix B.
4.3 Results
Figure 1 (right) shows the results of Experiment 2. Similar to the results of Experiment 1, the results on Pereira fMRI also show that as parameter count (model size) increases, the regression model log likelihood decreases significantly ($p$ < 0.05). The replication of the inverse scaling phenomenon on a different dataset suggests that this phenomenon is not dataset-specific. [^8] For both corpora, in addition to running experiments using LME models, per reviewers’ request, we also ran linear regression for both datasets (1) using one data point per sentence and (2) averaging across participants. For Pereira fMRI, since the response data is already per-sentence BOLD signals, for each sentence, we averaged across the BOLD signals across all participants. Similar to our method described in Experiment 2, which was based on the methodology in described Hosseini et al. (2024), we used the last-word surprisal estimate as the representation of each data point. We then performed linear regression modeling using six-fold cross-validation. The test fold sizes range from 87 to 122 data points and the training fold sizes range from 505 to 540 data points. For Natural Stories fMRI, since the response data is the time series of BOLD signals, we do not have direct information about sentence segmentation. In order to obtain an estimate of the sentence-level BOLD signals and surprisal estimates, in general, we followed Feghhi et al. (2024) which used an fMRI dataset collected at a two-second time interval (Blank et al., 2014), similar to Natural Stories fMRI. We mapped each response data point to the word (along with its surprisal estimate) that is closest to four seconds before that data point. Note that for this analysis, we did not apply HRF convolution to the predictors. This allowed us to obtain an estimate of sentence segmentation points since we have the sentence id information of each word. To get the sentence-level BOLD signals, we averaged all BOLD signals that were mapped to the same sentence id. As for sentence-level surprisal estimates, it was not possible to use last-word surprisal since the BOLD signals are not sentence-aligned, so for each sentence, we averaged all surprisal estimates that were within that sentence and were mapped to some BOLD signals. Once we obtained the sentence-level BOLD signals for each participant and surprisal estimates, we averaged across participants’ responses. Since the number of data points were reduced to a small amount after obtaining the sentence-level responses/surprisal estimates and averaging across subjects, we performed linear regression modeling using eight-fold cross-validation. For each iteration, we used the data of one story as the test data and the remaining seven stories as the training data. The test fold sizes range from 210 to 390 data points and training fold sizes range from 1992 to 2172 data points. For both corpora in this analysis, we report the mean squared errors averaged across each iteration. In this analyses, Natural Stories fMRI still shows a significant inverse scaling (p ¡ 0.05) while Pereira fMRI shows numerically but not statistically significant inverse scaling. We speculate that the reason why statistical significance was not observed for inverse scaling on Pereira fMRI under this experimental setting might be because the BOLD signals used for regression modeling were averaged across subjects, which could potentially make the data somewhat sparser and therefore underpowered.
5 Conclusion
Recent studies (Oh et al., 2022; Oh and Schuler, 2023; Oh et al., 2024) have observed a robust negative relationship between Transformer-based LMs’ quality and the predictive power of surprisal estimates from those LMs on latency data. This work generalizes that result by examining the predictive power of surprisal estimates from Transformer-based models on brain imaging data. The predictive power of surprisal estimates from 17 pre-trained LMs across three LM families was evaluated on two fMRI datasets. Results show that the inverse scaling between model size and model fit still obtains on brain imaging data. The inverse relationship between the predictive power of surprisal estimates from those LMs and parameter count (model size) on brain imaging data is not as strict as that on latency-based data, but this may be due to the fact that brain imaging data is much noisier than latency-based data. Nevertheless, the observation of the inverse scaling phenomenon on two distinct brain imaging datasets indicates that this trend is not specific to latency-based measures, suggesting that brain imaging data do not behave substantially differently from latency data. We speculate the deviation of larger LMs from brain imaging data is due to the same reason for reading times (Oh and Schuler, 2023). This additional piece of evidence reinforces the observation of inverse scaling phenomenon on latency data (Oh et al., 2022; Oh and Schuler, 2023), suggesting that smaller models could potentially be helpful in providing researchers with insights into the appropriate use of LMs in understanding sentence processing.
Limitations
This study attempts to examine the predictive power of surprisal estimates from Transformer-based models on fMRI data. Large language models evaluated in this work are trained on English text and the datasets evaluated are collected from English speakers. Therefore, these findings may or may not be replicated in other languages.
Ethics Statement
The fMRI datasets used in this study are from published work (Shain et al., 2020; Pereira et al., 2018). Details regarding the data collection, validation and other relevant procedures are described in respective publications. Since this work focuses on evaluating the predictive power of large language models on fMRI data, the potential risks and harmful impacts posed by this study on the society appear to be minimal.
Acknowledgments
We thank the anonymous reviewers for their helpful feedback. This work was supported by the National Science Foundation grant #2313140. All views expressed are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Appendix A Hyperparameters of Models Examined in this Study
Model Variant #L #H $d_{model}$ Parameters
GPT-2 Small 12 12 768 $\sim$124M
GPT-2 Medium 24 16 1024 $\sim$355M
GPT-2 Large 36 20 1280 $\sim$774M
GPT2- XL 48 25 1600 $\sim$1.6B
GPT-Neo 125M 12 12 768 $\sim$125M
GPT-Neo 1.3B 24 16 2048 $\sim$1.3B
GPT-Neo 2.7B 32 20 2560 $\sim$2.7B
GPT-J 6B 28 16 4096 $\sim$6B
GPT-NeoX 20B 44 64 6144 $\sim$20B
OPT 125M 12 12 768 $\sim$125M
OPT 350M 24 16 1024 $\sim$350M
OPT 1.3B 24 32 2048 $\sim$1.3B
OPT 2.7B 32 32 2560 $\sim$2.7B
OPT 6.7B 32 32 4096 $\sim$6.7B
OPT 13B 40 40 5120 $\sim$13B
OPT 30B 48 56 7168 $\sim$30B
OPT 66B 64 72 9216 $\sim$66B
Table 1: Hyperparameters of the models examined in this study. #L refers to the number of layers of that model; #H refers to the number of attention heads each layer has; $d_{model}$ refers to the embedding size of the model.
Appendix B Maximal LME models
During LME model fitting, we followed Barr et al. (2013) and started with the maximal LME models. We eliminated one random effect at a time until the LME models converge. To ensure that the regression includes both by-subject and by-item random effects, we removed random effects iteratively from the by-subject and by-item effects. That is, if the LME model does not converge, we first removed a random effect from the by-item group; if the LME model still does not converge, we then move on to removing a random effect from the by-subject group, and so on.
Following is the original design of the LME model for Natural Stories fMRI, in which the eliminated random effects are bolded.
BOLD ˜ z.(sentpos) + z.(wlen) + z.(surp) + (1 + z.(sentpos) + z.(wlen) + z.(surp) | subject) + (1 + z.(sentpos) + z.(wlen) + z.(surp) | storyid)
As for Pereira fMRI, the original design of the LME model is the following, with the eliminated random effect bolded:
BOLD ˜ z.(sentlen) + z.(passpos)+ z.(lastwordsurp) + (1 + z.(sentlen) + z.(passpos) + z.(lastwordsurp) | subject) + (1 + z.(sentlen) + z.(passpos) + z.(lastwordsurp) | sentid)